a projectile is thrown at an angle 37degree from the vertical the angle of elevation of the highest point of the projectile from point of projection is

Grade:12th pass

1 Answers

Arun
4 years ago
Dear studentIn projectile2H/R = 1/2 tan Ѳ = ½ tan 37 = 3/8
 
Hope it helps
 
Regards
Arun (askIITians forum expert)